DM Monthly Status Report for February 2019

The DM monthly status report covering February activities has been posted to DocuShare, collection-742. For convenience, the High-level Summary is pasted below. Direct link to the full report (pdf):

High-level Summary

DM held the first virtual DMLT F2F meeting Feb 26-28; many technical discussions were held and all agreed this was a good way to meet though we should still have physical meetings at times. Some DM team members participated in the Kavli funded Petabytes to Science Workshop in Las Vegas, NV and worked on recommendations for the decadal survey.

DM supported the DESC winter collaboration meeting, DM liaisons (Guy, Lupton) attended and discussed many topics concerning DM & DESC collaborative work. Simon Krughoff ran a demo of the LSST science platform.

The call for letters of intent for community brokers went out and preparations started for the broker workshop, including issuing of an Alerts Key numbers technote, DMTN-102.

Major upgrades to the “Match Pessimistic B” astrometric matcher, which have caused substantial improvements to the astrometric performance metrics, were made.

Support for color terms were added to the Jointcal calibration tool. We now have a task for injecting simulated sources into visit images. An extensive analysis of fitting complex galaxy models to the COSMOS field was completed. Further work will be required to develop metrics which monitor the quality of the fitting and figure out computational cost. We finally merged the obs_lsst camera support package for LSST to the codebase and unblocked Release 17 of the pipelines.

An impressive “generation 3” middleware demo was given early in Feb involving NCSA, Architecture, and Pipelines. This was a demonstration of integration of the Pegasus Workflow Management System with the current release of pipeline tasks and middleware.

A draft of LSE-400 defining the Header Service and the metadata it is to produce was completed and submitted to LSST Systems Engineering. A draft of LSE-349 defining the transformation between Camera Engineering coordinates and Camera Data Visualization coordinates was issued for discussion.

Authentication for the LSST Science Platform was integrated with the Portal Aspect and the WebDAV file management web service. A TAP query service backed by Qserv was deployed in the lsst-lsp-int and lsst-lsp-stable science platform instances.

The first version of the Observatory Operations Data System (OODS) software, which will provide low-latency access to images, other files, and metadata for use by Observatory systems and the Commissioning Cluster Science Platform instance was completed.

NCSA hosted and participated in a workshop that included installation of new Camera Data Acquisition (DAQ) hardware on the L1 Complete Test Stand at NCSA as well as a tutorial for the new software interface.

The Santiago ring and La Serena - Santiago primary and backup links are fully operational. AURA is now using 4x10G. We completed the conduits, trays, and network cabling for the Camera rooms. The GPS/Time Server cable is installed and the GPS is getting time from the server. The Weather tower and All-Sky Camera fibers are spliced inside the cabinets and connected to the computer room. Some work remains on the network switch and power for those as well as for the Differential Image Motion Monitor (DIMM) which will be delivered in March.